# Trading robustness: a scenario-free approach to robust Multi-Criteria Optimization for Treatment Planning

### v1

@@ 1 Introduction
The primary objective of radiotherapy treatment planning is to achieve an optimal balance between clinical goals, including adequate coverage of the target volume and minimal exposure of the surrounding organs at risk (OARs). As clinicians must navigate trade-offs between these competing clinical goals, the treatment planning problem inherently represents a multi-criteria decision making problem. In treatment planning, this problem has been traditionally approached by representing the trade-offs as a weighted sum of objectives to be optimized with scalar optimization techniques. In practice, this often results in a tedious iterative tuning process to find clinically acceptable plans whose quality further depends on the planners’ experience [1, 2]. To systematize the navigation of the planning trade-offs, more sophisticated multi-criteria optimization (MCO) approaches have been introduced into clinical workflows [3, 4, 5, 6, 7, 8, 9]. These methods offer advantages over traditional weighted-sum approaches by enabling consistent and automated navigation of a priori goals or by allowing a fast and interactive a posteriori navigation of available solutions [7, 5]. Common examples for a priori methods are hierarchical or lexicographic optimization[3, 10], while the most common a posteriori method relies on Pareto surface approximation [4, 8].
Beyond the challenges of decision-making, the treatment plan quality can be significantly affected by uncertainties such as anatomical variations, patient setup or range errors[11, 12, 13, 14]. According to the specific radiation modality and treatment site, robust optimization techniques can effectively reduce the required margin size and the integral dose to the patient [15, 16].
However, traditional robust optimization approaches involve evaluationg multiple error scenarios during both dose calculation and optimization which makes them computationally expensive. At the same time, standard MCO methods, typically require the sequential generation of multiple plans, exacerbating the computational complexity of the problem and scaling the optimization time and memory footprint to unfeasibility.
Conventional robust optimization algorithms rely on well defined mathematical metrics such as max value operators, for min-max optimization approaches or expected value operators for probabilistic planning [16, 15]. Despite their rigorous formulation, predicting the impact of these operators on the final plan remains challenging, and their effectiveness closely relates to the accuracy of the uncertainty model employed. These operators influence the problem formulation in binary manner, i.e once applied, they are enforced fully on all solutions.
@@
Several approaches tackled the first practical limitation in both optimization and robustness analysis. Examples include Polynomial Chaos Expansion [17, 18] and Analytical Probabilistic Models (APM) [19, 20], which utilize analytical methods to propagate uncertainty through dose calculation. These techniques efficiently estimate the first moments of the dose for each voxel, and eliminate the explicit calculation of all error scenarios during optimization. However, they often still entail significant computational costs associated with constructing the analytical model itself.
A solution to overcome the second limitation, and enable robust MCO, was proposed in literature [21] where traditional robust operators are applied to selected objectives. A combination of robust and non robust objectives builds the Pareto front approximation allowing the control of plan robustness through trade-off between these objectives. This approach offers a viable solution to the conceptual limitations of robust optimization, but still bares the computational burden of explicit individual scenario evaluation during optimization.
@@
Variance can be estimated during optimization through the explict evaluation of all error scenarios, nonetheless baring again the same computational cost. Alternatively, the recently proposed scenario-free approach to robust optimization enables the natural and efficient integration of mean variance reduction objectives [22]. This efficiency arises from precalculation of probabilistic quantities over the set of error scenarios, avoiding explicit evaluation of each scenario during optimization.
@@
In this work we propose and investigate the application of the scenario-free approach to two MCO techniques, namely a Lexicographic Ordering (LO) and a Pareto front approximation and navigation approach. The LO strategy is a so-called (2pc) approach [7, 10] while the Pareto front approximation is obtained through a sandwiching algorithm [23, 24, 25]. Several optimization configurations are explored to investigate the impact of variance reduction as a decision making criterion for the different strategies.
@@ 2.1.1 A posteriori approach: Pareto Front navigation
This “trial-and-error” process of tuning weights, waiting for the next solution, and repeating with an adapted set of weights , may practically be replaced by an automated strategy to generate plans with different sets of weights . The method of choosing can be formalized, for example using sandwiching algorithms, to adequately cover the Pareto front. Combined with a navigation technique to interpolate between generated points, the clinical decision is then performed a-posteriori. The planner can explore the tradeoffs between objectives and their impact on the plan metrics by intuitive graphical user interfaces.
@@ 2.2 The scenario-free approximation to robust optimization
As an alternative approach to robust optimization, the scenario-free approach [22] enables efficient evaluation of the expected dose distribution and mean variance during optimization. These quantities are estimated from precomputed probabilistic quantities through:
@@ 2.4.1 Objective functions and clinical goals
The variance reduction objective was formulated as minimization of the variance term in eq. 5. The equivalent uniform dose (EUD) objective depends on the parameter as defined in table 1. This parameter encodes the volume effect of the specific organ and for parallel OARs, values of are usually applied[26, 27]. An higher value of increases the penalty and the attention of the optimizer for higher doses. In this work a value of was chosen. This choice allows to maintain a clear clinical interpretation of the goal value associated to this objective while inducing an stronger reduction of higher doses in the lung. This choice additionally ensures the convexity of the EUD function [28].
@@
Table 2 collects the priority lists designed for the strategies (Strategy 1-4). Variance reduction for the target structure is prioritized differently for each strategy described in section 2.3.1, balancing its relative importance against the dosimetric objectives for the OARs. For all strategies, an additional minimum dose constraint is applied to the target at of the prescribed dose.
@@
For the Pareto Front approximation, multiple objective functions for the same structure are combined and treated as single objectives (table 1). For the OARs, this includes a least-squares deviation on the EUD plus a squared overdose objective on dose, forming a single objective. A variance minimization term is applied on the target. Additional minimum and maximum dose constraints are applied to the target to ensure coverage and avoid hot spots. The minimum and maximum target doses are set to and .
@@
2.4.2 Patient datasets
In this work, a lung cancer patient dataset was selected from the cancer imaging archive [29]. The dataset consists of a 4D-CT scan with phases representing the breathing motion. This choice allows for the separate investigation of the different considered sources of uncertainty, particularly of setup and range uncertainties as well as anatomical motion.
2.5 Robustness Analysis
2.6 Algorithm implementation
Both the scenario-free approach and the LO approach were implemented in the open-source treatment planning toolkit matRad [30]. matRad served as the development environment for individual scenario dose calculations, accumulation of probabilistic quantities, and the final robustness analysis performed on all obtained solutions.
3 Results
3.1 LO approach
3.1.1 Strategy 1: Algorithm validation
Finally, the relative optimization times for all steps range between 1.5 and 1.6. This indicates that the s-f approach is approximately 1.5 times slower than the nominal optimization. Such overhead is introduced by the additional computation of the variance operator. Variance estimation is performed at each optimization step as a constraint value on the variance is always applied during the second phase of the 2pc approach.
3.1.2 Strategy 2: Prioritization of dosimetric objectives
3.1.3 Strategy 3: Prioritization of Variance reduction objectives
The third prioritization strategy places variance reduction within the target volume as the primary objective. Figure 1, Strategy 3 presents the DVH and SDVH analyses for the robust plan at the last steps of the optimization process. A highly robust dose distribution is achieved for the target as early as the first step of optimization and minimal variance is maintained throughout the remaining optimization steps.
The dose distribution reported for this strategy shows a lower target conformity and higher OARs doses compared to the results presented in section (3.1.1) and (3.1.2). At the same time the SD distributions and SDVH curves highlight how the standard deviation is highly reduced in the target compared to the previous approaches.
3.1.4 Strategy 4: 4D robustness
The fourth optimization strategy was designed to address setup and range (3D) uncertainties separately from breathing motion (4D) uncertainties. Figure 2 illustrates the robustness analysis associated with this strategy, focusing on the variance reduction achieved in Steps 5 and 6. In Step 5, 4D variance reduction is performed, and the robustness analysis was conducted on both the 3D-only and the 4D-only scenario datasets. For both robustness analyses, the figure also reports the SDVH for the target structure at the previous dosimetric-only step (Step 4) for direct comparison.
Figure 2 also presents the DVH and SDVH analysis for the final optimization step (Step 6), in which the 3D variance reduction is performed. At this stage, the robustness analysis was performed on the combined dataset of 3D and 4D scenarios, thereby accounting for the total uncertainty arising from both sources. For comparison, the SDVH plot also includes the curves obtained for the target, on the same dataset, in Steps 4 and 5, corresponding to dosimetric-only optimization and 4D-only variance reduction, respectively. Across these steps, the SDVH curve progressively shifts to the left, reflecting the reduction in uncertainty.
3.2 Pareto Front approximation
These trade-offs can be explored interactively through interpolation of the Pareto surface, allowing for dynamic assessment of alternative solutions. This functionality is enabled by a graphical decision-support tool developed by the Fraunhofer ITWM . The tool includes real-time visual updates of objective projections and DVH comparisons.
4 Discussion
The scenario-free (s-f) approach enables the efficient evaluation of both the expected dose distribution and the mean variance term during optimization. While these quantities could also be computed using traditional robust optimization methods and evaluating all the included scenarios, such approach is typically hindered by high computational demands and memory limitations. In MCO workflows, where multiple sequential optimizations are performed, these computational costs accumulate, making the use of traditional robust optimization techniques impractical. In contrast, the s-f approach offers a fast alternative for achieving robust optimization within MCO, requiring optimization times comparable to those of nominal, non-robust approaches.
Traditional approaches allow only a binary application of robust optimization, meaning that the robustness component is either fully included in the problem formulation or entirely excluded. The choice of the robustness operator or margin dictates the nature of the optimized plan, which may become too optimistic or too conservative depending on the settings [15, 33, 34]. Because of this rigidity, the influence of robust optimization on the final dose distribution can only be adjusted by adapting the uncertainty source [35, 36, 37], or, when possible, by adopting compromising strategies [38]. However, the selection of an uncertainty model should be guided by evidence and clinical practice [39, 40] whenever sufficient data is available.
In contrast, the scenario-free approach incorporates variance reduction as a criterion in the decision making process. This enables explicit control of plan robustness through variance reduction, which is not bounded to any a-priory choice of robustness operator. Moreover, if the underlying uncertainty model does not accurately represent the real, undetermined, probability distribution of the error source, the scenario-free approach still provides a consistent solution to practically manage the impact of such uncertainty on the final dose distribution.
4.1 Prioritization strategies for the LO approach
The lexicographic ordering framework enables the explicit assignment of priority levels to planning objectives, offering a clear and clinically interpretable structure. LO also relies on the assumption that all objectives can be ranked. The ranking priority may be clear for traditional dosimetric goals, when the DM relies on extensive planning experience, but is less clear for variance reduction objectives. In this work, we explored four prioritization strategies to examine how different objective orderings affect the resulting treatment plans.
The subsequent variance reduction step effectively reduces the variance within the targeted structure thus proving the efficacy of the variance reduction objective in limiting the impact of uncertainty while still maintaining the achieved dosimetric plan quality.
This prioritization strategy yields an exceptionally robust dose distribution for the target, which remains stable across all subsequent optimization steps. However, this robustness comes at the expense of target conformity and reduced sparing of nearby organs at risk, highlighting the trade-offs inherent in such prioritization.
Regardless of the specific prioritization strategy, each step in the LO sequence allows for a controlled relaxation of the previously optimized objectives. This is managed through the introduction of the slack variable , [7], which ensures numerical stability during the optimization process. No significant degradation in the quality of any previously optimized objective was observed across the analyzed cases, for any of the applied Strategies.
4.2 Pareto front approximation
This analysis focused on a three-dimensional objective space, allowing full visualization of the Pareto front and direct exploration of how variance reduction objectives affect the overall plan quality. The representation clearly demonstrates how robustness, especially in the target, trades off against sparing of OARs.
All sampled solutions lie on the Pareto front, guaranteeing Pareto efficiency and enabling the decision-maker to explore meaningful trade-offs between competing objectives. Combining this approach with the computational speed of the s-f framework allows efficient construction of the Pareto front approximation, effectively bringing robust Pareto front exploration within the reach of clinical use. The interactive decision-support tool further enhances this process by enabling real-time exploration of the Pareto surface, DVHs, and dose distributions, supporting informed clinical choices.
4.3 Summary and Outlook
Hybrid strategies that combine the strengths of both approaches are also possible. For instance, a lexicographic, clinically informed ordering could be applied to the dosimetric objectives, followed by a Pareto front approximation performed solely in the variance reduction space. This would allow robustness to be balanced across different structures, or different sources of uncertainty, without compromising initial dosimetric goals. Additionally, the dual nature of the problem – addressing both dose and variance – suggests the potential for a bi-level optimization formulation. Recent work demonstrated the effectiveness of bi-level approaches for dose and LET optimization, particularly when supported by an efficient approximation of the domination cone [41]. A similar formulation could be explored in the context of dose and variance optimization as well.
5 Conclusions
6 Aknowledgement
Funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), projects no. 443188743 and 457509854, is acknowledged.
7 Conflict of interest

### v2

@@ 1 Introduction
The primary objective of radiotherapy treatment planning is to achieve an optimal balance between clinical goals, including adequate coverage of the target volume and minimal exposure of the surrounding organs at risk (OARs). As clinicians must navigate trade-offs between these competing clinical goals, the treatment planning problem inherently represents a multi-criteria decision making problem. In treatment planning, this problem has been traditionally approached by representing the trade-offs as a weighted sum of objectives to be optimized with scalar optimization techniques. In practice, this often results in a tedious iterative tuning process to find clinically acceptable plans whose quality further depends on the planners’ experience [undef, undefa]. To systematize the navigation of the planning trade-offs, more sophisticated multi-criteria optimization (MCO) approaches have been introduced into clinical workflows [undefb, undefc, undefd, undefe, undeff, undefg, undefh]. These methods offer advantages over traditional weighted-sum approaches by enabling consistent and automated navigation of a priori goals or by allowing a fast and interactive a posteriori navigation of available solutions [undeff, undefd]. Common examples for a priori methods are hierarchical or lexicographic optimization[undefb, undefi], while the most common a posteriori method relies on Pareto surface approximation [undefc, undefg].
Beyond the challenges of decision-making, the treatment plan quality can be significantly affected by uncertainties such as anatomical variations, patient setup or range errors[undefj, undefk, undefl, undefm]. According to the specific radiation modality and treatment site, robust optimization techniques can effectively reduce the required margin size and the integral dose to the patient [undefn, undefo].
However, traditional robust optimization approaches involve evaluating multiple error scenarios during both dose calculation and optimization which makes them computationally expensive. At the same time, standard MCO methods, typically require the sequential generation of multiple plans, exacerbating the computational complexity of the problem and scaling the optimization time and memory footprint to unfeasibility.
Conventional robust optimization algorithms rely on well defined mathematical metrics such as max value operators, for min-max optimization approaches or expected value operators for probabilistic planning [undefo, undefn]. Despite their rigorous formulation, predicting the impact of these operators on the final plan remains challenging, and their effectiveness closely relates to the accuracy of the uncertainty model employed. These operators influence the problem formulation in binary manner, i.e once applied, they are enforced fully on all solutions.
The choice of the robustness operator or margin dictates the nature of the optimized plan, which may become too optimistic or too conservative depending on the settings [undefn, undefp, undefq]. Because of this rigidity, the influence of robust optimization on the final dose distribution can only be adjusted by adapting the model describing the uncertainty source and its parameters [undefr, undefs, undeft], or, when possible, by adopting balancing strategies [undefu].
@@
Several approaches tackled the first practical limitation in both optimization and robustness analysis. Examples include Polynomial Chaos Expansion [undefv, undefw] and Analytical Probabilistic Models (APM) [undefx, undefy], which utilize analytical methods to propagate uncertainty through dose calculation. These techniques efficiently estimate the first moments of the dose for each voxel, and eliminate the explicit calculation of all error scenarios during optimization. However, they often still entail significant computational costs associated with constructing the analytical model itself.
A solution to overcome the second limitation, and enable robust MCO, was proposed in literature [undefz] where traditional robust operators are applied to selected objectives. A combination of robust and non robust objectives builds the Pareto front approximation allowing the control of plan robustness through trade-off between these objectives. This approach offers a viable solution to the conceptual limitations of robust optimization, but still bares the computational burden of explicit individual scenario evaluation during optimization.
@@
Variance can be estimated during optimization through the explict evaluation of all error scenarios, nonetheless baring again the same computational cost. Alternatively, the recently proposed scenario-free approach to robust optimization enables the natural and efficient integration of mean variance reduction objectives [undefaa]. This efficiency arises from precalculation of probabilistic quantities over the set of error scenarios, avoiding explicit evaluation of each scenario during optimization.
@@
In this work we propose and investigate the application of the scenario-free approach to two MCO techniques, namely a Lexicographic Ordering (LO) and a Pareto front approximation and navigation approach. The LO strategy is a so-called two-phase -constrained (2pc) approach [undeff, undefi] while the Pareto front approximation is obtained through a sandwiching algorithm [undefab, undefac, undefad]. Several optimization configurations are explored to investigate the impact of variance reduction as a decision making criterion for the different strategies.
@@ 2.1.1 A posteriori approach: Pareto Front navigation
This “trial-and-error” process of tuning weights, waiting for the next solution, and repeating with an adapted set of weights , may practically be replaced by an automated strategy to generate plans with different sets of weights . The method of choosing can be formalized, for example using sandwiching algorithms, to adequately cover the Pareto front. Combined with a navigation technique to interpolate between generated points, the clinical decision is then performed a posteriori. The planner can explore the tradeoffs between objectives and their impact on the plan metrics by intuitive graphical user interfaces.
@@ 2.2 The scenario-free approximation to robust optimization
As an alternative approach to robust optimization, the scenario-free approach [undefaa] enables efficient evaluation of the expected dose distribution and mean variance during optimization. These quantities are estimated from precomputed probabilistic quantities through:
@@ 2.4.1 Objective functions and clinical goals
The variance reduction objective was formulated as minimization of the variance term in eq. 5. The equivalent uniform dose (EUD) objective depends on the parameter as defined in table 1. This parameter encodes the volume effect of the specific organ and for parallel OARs, values of are usually applied[undefae, undefaf]. An higher value of increases the penalty and the attention of the optimizer for higher doses. In this work a value of was chosen. This choice allows to maintain a clear clinical interpretation of the goal value associated to this objective while inducing a stronger reduction of higher doses in the lung. This choice additionally ensures the convexity of the EUD function [undefag].
@@
Table 2 in section 8 collects the priority lists designed for the strategies (Strategy 1-4). Variance reduction for the target structure is prioritized differently for each strategy described in section 2.3.1, balancing its relative importance against the dosimetric objectives for the OARs. For all strategies, an additional minimum dose constraint is applied to the target at of the prescribed dose. The definition of target structure differs between the scenario-free and the nominal case. For the nominal case this is set to the PTV, while for the scenario-free approach it is always set to the CTV.
@@
For the Pareto Front approximation, multiple objective functions for the same structure are combined and treated as single objectives (table 1). For the OARs, this includes a least-squares deviation on the EUD plus a squared overdose objective on dose, forming a single objective. A variance minimization term is applied on the CTV. Additional minimum and maximum dose constraints are applied to the CTV to ensure coverage and avoid hot spots. The minimum and maximum target doses are set to and .
2.4.2 Patient datasets
In this work, a lung cancer patient dataset was selected from the cancer imaging archive [undefah]. The dataset consists of a 4D-CT scan with phases representing the breathing motion. This choice allows for the separate investigation of the different considered sources of uncertainty, particularly of setup and range uncertainties as well as anatomical motion.
2.5 Robustness Analysis
These metrics provide a quantitative measure of uncertainty and a measurable impact of the optimization strategy. However, plan robustness is conventionally assessed by testing selected metrics against predetermined acceptance criteria, and a plan is considered robust when the selected metric reaches the desired quality for a predefined percentage of scenarios.
When variance is minimized with the scenario-free approach, there is no rigorous mathematical way of deriving such confidence intervals from the scalar mean variance value estimated. Testing of the desired metrics against the scenario set needs to be performed for the selected solution, a posteriori and through robustness analysis.
Previous work [undefaa] validated the proposed scenario-free approach against scenario-based probabilistic planning however, the proposed MCO integration shifts the focus toward decision making and robustness as a criterion. Conventional worst-case optimization serves thus as a more appropriate reference and is addressed by initial analysis.
The link between conventional robustness metrics and variance is investigated with the following strategy: a plan is initially optimized with a conventional composite-worst case optimization approach where the scalarized cost function with the maximum value among the error scenarios is minimized. A total of setup and range error worst-case scenarios in addition to the nominal scenario are collected from the same probability distributions described in section 2.4.2 according to Korevaar et al.[undefai]. The worst-case errors are sampled at covering of the uncertainty space. The plan is optimized such that a CTV coverage of is achieved for at least of the worst-case scenarios.
Three scenario-free plans are optimized using the same set of worst-case scenarios to compute the required probabilistic quantities. The objectives for the scenario-free plan are equivalent to the ones used for the worst-case optimization, with the only difference of being applied to the expected dose distribution as opposed to the individual scenario distributions. Additional variance reduction objectives are applied to the CTV, lung and heart. The relative importance of the variance reduction objectives is set and progressively relaxed for the three plans to: a) match the DVHs and SDVHs of the worst-case plan, thus obtaining an equivalent plan meeting the same acceptance criteria, b) obtain a less conservative result, closer to probabilistic optimization and still meeting the robustness criteria, c) further prioritize OAR sparing against target coverage but failing the acceptance criteria. The three s-f plans showcase how different priorities set for the variance reduction objectives can reproduce equivalent conventional results while providing the degree of freedom to explore solutions within the robustness acceptance range.
The robustness analysis for these plans and the following MCO strategies is performed on the corresponding, complete probabilistic scenario set used for optimization. The reported analysis is thus a probabilistic evaluation of the plans. The chosen robustness criteria addresses target coverage as the number of scenarios for which at least of the CTV volume receives at least of the dose prescription (). The acceptance threshold for this analysis is arbitrarily set at of the scenarios as well.
The passing criteria analysis is reported for the LO strategies alone, while the Pareto front analysis only focuses on the exploratory aspects of the proposed method with relative reduction of variance.
2.6 Algorithm implementation
Both the scenario-free approach and the LO approach were implemented in the open-source treatment planning toolkit matRad [undefaj]. matRad served as the development environment for individual scenario dose calculations, accumulation of probabilistic quantities, and the final robustness analysis performed on all obtained solutions.
3 Results
3.1 Conventional robustness reference
A plan is initially optimized using the conventional worst-case optimization and the worst-case scenarios set and compared against three s-f plans obtained relaxing the importance weight for the variance reduction objective. Robustness analysis is performed on the randomly sampled set of 3D scenarios.
Figure 1 reports the robustness analysis performed on all plans. The worst-case solution and the s-f solution with tight variance reduction result in equivalent and overlapping DVHs, DVH bands and SDVHs for the CTV structure and the lung, practically resulting in equivalent plans fulfilling the same robustness acceptance criteria. The percentage of probabilistic scenarios meeting the robustness criteria () is for both plans.
The same analysis obtained for the mild- and weak-variance reduction scenario-free plans shows decreased target coverage and increased variability over the scenarios. Accordingly, the DVHs for the lung follow the opposite trend, showing higher sparing of the OAR with the increasing CTV variance.
The robustness criteria is met by and of the probabilistic scenarios for the mild- and weak-variance reduction plans, respectively. The mild-variance plan still satisfies the acceptance threshold set at while the weak-variance reduction plan does not.
3.2 LO approach
3.2.1 Strategy 1: Algorithm validation
The reported robustness analysis was performed with all randomly sampled error scenarios used for the precalculation of the probabilistic quantities and is thus a probabilistic analysis. The percentage of scenarios meeting the robustness condition is of for the margin based and for the s-f robust plan. With the acceptance rate set at the margin-based approach would thus not be accepted, as opposed to the s-f plan.
Finally, the relative optimization times for the scenario-free approach range between 1.5 and 1.6 for all steps of the LO optimization sequence, meaning the scenario-free optimization is approximately 1.5 slower compared to the nominal optimization. Such overhead is introduced by the additional computation of the variance objective at each iteration of the optimizer. Total variance calculation is performed at each optimization step due to the total variance constraint applied in the second phase of the 2pc approach. The rTPI refers to the optimization step alone and does not include the dose calculation step, which still needs to be performed for all scenarios.
3.2.2 Strategy 2: Prioritization of dosimetric objectives
The percentage of scenarios meeting the robustness criteria moved from to for Step and step respectively denoting the relation between variance reduction and improvement of conventional robustness metrics. The passing rate for the final plan would be close but not sufficient to meet the acceptance threshold set at .
3.2.3 Strategy 3: Prioritization of Variance reduction objectives
The third prioritization strategy places variance reduction within the CTV volume as the primary objective. Figure 2, Strategy 3 presents the DVH and SDVH analyses for the robust plan at the last steps of the optimization process. A highly robust dose distribution is achieved for the target as early as the first step of optimization and minimal variance is maintained throughout the remaining optimization steps.
The dose distribution reported for this strategy shows a lower target conformity and higher OARs doses compared to the results presented in section (3.2.1) and (3.2.2). At the same time the SD distributions and SDVH curves highlight how the standard deviation is highly reduced in the CTV compared to the previous approaches.
of the scenarios meet the robustness criterion already at the first step of the optimization and this is maintained until the last optimization step.
3.2.4 Strategy 4: 4D robustness
The fourth optimization strategy was designed to address setup and range (3D) uncertainties separately from breathing motion (4D) uncertainties. Figure 3 illustrates the robustness analysis associated with this strategy, focusing on the variance reduction achieved in Steps 5 and 6. In Step 5, 4D variance reduction is performed, and the robustness analysis was conducted on both the 3D-only and the 4D-only scenario datasets. For both robustness analyses, the figure also reports the SDVH for the CTV structure at the previous dosimetric-only step (Step 4) for direct comparison.
Figure 3 also presents the DVH and SDVH analysis for the final optimization step (Step 6), in which the 3D variance reduction is performed. At this stage, the robustness analysis was performed on the combined dataset of 3D and 4D scenarios, thereby accounting for the total uncertainty arising from both sources. For comparison, the SDVH plot also includes the curves obtained for the CTV, on the same dataset, in Steps 4 and 5, corresponding to dosimetric-only optimization and 4D-only variance reduction, respectively. Across these steps, the SDVH curve progressively shifts to the left, reflecting the reduction in uncertainty.
3.3 Pareto Front approximation
These trade-offs can be explored interactively through interpolation of the Pareto surface, allowing for dynamic assessment of alternative solutions. This functionality is enabled by a graphical decision-support tool developed by the Fraunhofer ITWM . The tool includes real-time visual updates of objective projections and DVH comparisons. An example of the use and functionality of this tool is available to the reader as additional material.
4 Discussion
The scenario-free approach incorporates variance reduction as a criterion in the decision making process. This enables explicit control of plan robustness through variance reduction, which is not bounded to any a-priori choice of robustness operator. Moreover, if the underlying uncertainty model does not accurately represent the real, undetermined, probability distribution of the error source, the scenario-free approach still provides a consistent solution to practically manage the impact of such uncertainty on the final dose distribution.
Equivalency between scenario-free optimization and conventional robust optimization approaches may still be achieved by appropriate choice of the problem formulation. Previous work focused on the benchmarking of this approach against probabilistic optimization [undefaa] and the analysis reported in section 3.1 provides a further comparison to worst-case optimization. This serves as an indicator that solutions equivalent to conventional robust optimization approaches are feasible for the s-f optimization as well. The range of equivalently robust solutions can be explored with the s-f approach accessing direct control on the mean variance of different structures the potentiality of which well suits a MCO approach.
There is no rigorous mathematical formulation allowing to derive acceptance criteria from non-local, mean variance values alone. Proper plan robustness can thus only be determined through a posteriori robustness analysis on a selected set of error scenarios.
However, it may be possible that robustness levels across patients remain comparable when similar goals on total variance are met. This would allow to determine reference or calibration points for future plans, facilitating more rigorous probabilistic planning with evidence-based upper bounds on variance ensuring acceptance in probabilistic evaluation after optimization. The analysis in section 3.1 underlines this correlation between variance and conventional acceptance criteria, which is further supplemented by the remaining analyses of DVHs and plan metrics.
Further, the ability to trade robustness a posteriori by navigating the Pareto front could actively help planners achieve the robustness goals. [undefam] highlights the importance of performing probabilistic evaluation after planning to establish the compliance of the solution with the set robustness criteria.
The s-f approach applied in a MCO framework instead would provide, if not directly the value of variance to meet the criteria, at least the objective direction in which to explore the space to improve the plan metrics.
4.1 Prioritization strategies for the LO approach
The subsequent variance reduction step effectively reduces the variance within the targeted structure thus proving the efficacy of the variance reduction objective in limiting the impact of uncertainty while still maintaining the achieved dosimetric plan quality. The robustness acceptance criteria set at increased with the additional variance reduction step underlying the correlation between the value of such objective and the corresponding robustness metrics. For this particular case the acceptance threshold set at was not met (only was reached), meaning a more robust plan can be achieved increasing the variance reduction priority or relaxing the goal values for the OAR objectives with higher priority.
This prioritization strategy yields a dose distribution for the CTV that fulfills the robustness criteria already at the first Step of the LO sequence and remains stable across all subsequent optimization steps. However, this robustness comes at the expense of target conformity and reduced sparing of nearby organs at risk, highlighting the trade-offs inherent in such prioritization.
Regardless of the specific prioritization strategy, each step in the LO sequence allows for a controlled relaxation of the previously optimized objectives. This is managed through the introduction of the slack variable , [undeff], which ensures numerical stability during the optimization process. No significant degradation in the quality of any previously optimized objective was observed across the analyzed cases, for any of the applied Strategies.
4.2 Pareto front approximation
The Pareto front representation and the reported analysis demonstrates how variance reduction, especially in the target, trades off against sparing of OARs.
The analysis on the acceptance criteria is not reported for these solutions as the focus is posed on the opportunity offered by the method to explore diverse solutions. The correspondence with conventional robust optimization and the relation between variance reduction and passing criteria is already explored for the previous configurations and can easily be translated to this approach as well.
Combining this approach with the computational speed of the s-f framework allows efficient construction of the Pareto front approximation, effectively bringing robust Pareto front exploration within the reach of clinical use. The interactive decision-support tool further enhances this process by enabling real-time exploration of the Pareto surface, DVHs, and dose distributions, supporting informed clinical choices.
4.3 Summary and Outlook
Hybrid strategies that combine the strengths of both approaches are also possible. For instance, a lexicographic, clinically informed ordering could be applied to the dosimetric objectives, followed by a Pareto front approximation performed solely in the variance reduction space. This would allow robustness to be balanced across different structures, or different sources of uncertainty, without compromising initial dosimetric goals. Additionally, the dual nature of the problem – addressing both dose and variance – suggests the potential for a bi-level optimization formulation. Recent work demonstrated the effectiveness of bi-level approaches for dose and LET optimization, particularly when supported by an efficient approximation of the domination cone [undefan]. A similar formulation could be explored in the context of dose and variance optimization as well.
5 Conclusions
6 Aknowledgement
NW acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), project no. 443188743. TB and NW acknowledge funding by the Cooperation Program in Cancer Research of the German Cancer Research Center (DKFZ) and Israel’s Ministry of Innovation, Science and Technology (MOST), project no. Ca216.
7 Conflict of interest
8 Appendix